# How Grounded is Wikipedia? A Study on Structured Evidential Support

### v1

@@ 3.2 Evidence Retrieval
    LEAD: Retrieve the body evidence sentences for a given lead claim
@@
    BODY: Retrieve the evidence sentences from a single cited source for a given body claim
@@
For LEAD and BODY, we report recall@{5,10} and NDCG@5 results on the PeopleProfiles test set using several widely used retrieval models: BM25 (Robertson et al., 1995), ColBERTv2 (Khattab and Zaharia, 2020; Santhanam et al., 2022), and Stella-v5 1.5B (Zhang et al., 2024). Since ENTITY requires retrieval of many more sentences, we instead report recall@100, as well as NDCG@{5,100}.
Main Results
Table 2 reports the main results for all three models on all three tasks. We consistently obtain our best results with ColBERTv2, which shows 2+ point gains across metrics on LEAD and ENTITY, and 6+ point gains on BODY relative to the next best method.
Evidence retrieval difficulty increases with query scope.
@@
We observe wide variability in the difficulty of different tasks, with highest NDCG@5 scores on BODY, followed by LEAD and then by ENTITY. Intriguingly, this ranking tracks the granularity of claims/queries, where body claims (BODY) tend to provide the most detailed information, lead claims (LEAD) present key high-level facts, and entity-level queries (ENTITY) represent a limiting case—seeking generic biographical information. Intuitively, highly specific body claims likely bear greater lexical and semantic similarity to their supporting material than the higher-level claims of leads or the entity-level queries do to theirs.
Task #Sents N@5 R@5 R@10
LEAD 1 88.20 88.90 72.39
2 57.29 68.42 52.77
3 33.72 46.83 32.13
BODY 1 85.71 91.93 75.78
2 66.13 79.68 59.71
3 48.28 66.01 45.98
@@
Table 3: Retrieval results for ColBERTv2 broken down by number of evidence sentences. Retrieval performance drops sharply as # evidence sentences increases.
Evidence retrieval difficulty increases with evidence complexity.
Table 3 presents retrieval results on LEAD and BODY broken down by number of annotated evidence sentences. Whereas retrieval performance is strong for single-sentence evidence, we observe double-digit drops in moving to 2- and 3-sentence evidence sets. This may be explained by the fact that evidential support is often compositional, requiring integration of independently non- or weakly supporting pieces of evidence via inference rules. Addressing this by indexing larger passages, though tempting, would severely curtail the ability to localize the evidence: the average distance between evidence sentences for dev set body claims with multi-sentence evidence sets is 8.7 sentences, expanding to 14.6 for lead claims. That we find evidence dispersed across multiple sentences even in Wikipedia indicates that the need to handle compositional evidence is not a niche concern.
Reasoning rerankers help.
The above observations suggest that effective retrieval of complex evidence demands more sophisticated methods than lexical or semantic similarity match. Recent work shows that reasoning-based rerankers achieve substantial gains on other complex retrieval tasks (Weller et al., 2025; Shao et al., 2025; Zhuang et al., 2025). Accordingly, we leverage Rank1-7B, a pointwise reranker based on Qwen 2.5 7B (Qwen et al., 2025) distilled from 635K reasoning traces for MS MARCO relevance judgments produced by DeepSeek R1 (Guo et al., 2025). We use Rank1 to rerank the top 10 evidence sentences from BM25 for LEAD and BODY and the top 100 for ENTITY. Results are in Table 2’s “+Rerank” rows, where we find large gains over first-stage retrieval across all metrics—pointing to a vital role for reasoning-based rerankers in complex evidence retrieval. Table 4 has fine-grained results.
@@ 4 Conclusion
We have presented a study of evidential support in Wikipedia and have introduced PeopleProfiles, a large new resource of fine-grained, multi-level support annotations on 1,500 Wikipedia articles and their cited sources. We have shown that: (1) a sizable fraction of Wikipedia lead claims are unsupported by their body sections, and Wikipedia body claims by their (publicly accessible) cited sources; (2) evidence retrieval for these claims grows much more challenging as query scope and evidence complexity increases; and (3) new reasoning-based rerankers enable much more effective retrieval of complex evidence. We release PeopleProfiles to aid future work on claim verification and on furthering understanding of Wikipedia as a knowledge base for modern NLP.
@@
We acknowledge several limitations of our work. First, PeopleProfiles focuses only on Wikipedia articles about people. We chose this focus because biographies tend to have a higher proportion of straightforward, uncontroversial facts relative to other domains (e.g. concepts or events). However, it is possible the support distributions or the difficulty of evidence retrieval for articles in these other domains could differ from what we observe here. Second, as we emphasize in the main text, our claims about evidential support extend only to publicly accessible, digital sources—those that a human or machine reader could readily use to verify an article’s claims. We therefore cannot make conclusions about support across all source types in Wikipedia. Third, we leverage GPT-4o-mini as an annotator to facilitate our large-scale bulk data collection. While the agreement we observe between this model and our human annotations is strong (§2), LLMs have their own response biases and may not be fully calibrated when providing scalar judgments (Lovering et al., 2024).
Finally, a major challenge for all NLP work on citation is that real-world citation practices routinely deviate from the highly idealized model, widely adopted by NLP practitioners, in which each citation-worthy sentence is expected to bear a citation. In practice, a citation may be given (e.g.) only once at the beginning of a paragraph if it is clear that the remainder of the paragraph concerns the same source. Alternatively, several citations may be provided together only in the final sentence of a paragraph. Such practices may obscure the mappings between claims and their source material, but they are the norm and it is incumbent on the NLP community to find better ways to accommodate them.
Since we restrict ourselves to body sentences that do bear citations, we cannot draw conclusions about the (likely substantial) number of sentences whose source may appear as a citation on some other sentence. However, we suspect our estimate of the proportion of supported body claims decomposed from citation-bearing sentences is an upper bound on the proportion we would obtain for all citation-worthy sentences.
@@
PeopleProfiles’s use of sources from MegaWika and our release of this data (via a CC-BY-4.0-SA license) is consistent with MegaWika’s own CC-BY-4.0-SA license. Our principle transformation of the original Wikipedia articles consists in the decomposition of claims, which is performed by an LLM (GPT-4o-mini), and which can result in subclaims that misrepresent the article’s original content and thus (potentially) facts about the subject. Although our claim decompositions are highly faithful to the original texts, users of PeopleProfiles should be aware of this possibility.
@@ A.1 Annotator Demographics
The first three authors of this work, all native English-speaking graduate or professional NLP researchers, conducted the human pilot annotations. These authors also jointly produced the annotation instructions beforehand. None was compensated beyond their co-authorship on this work.
@@ A.2 Claim Decomposition
We use the “DnD” decomposition and decontextualization method introduced by Wanner et al. (2024b), which uses an LLM prompt-based method for obtaining subclaim decompositions and the corresponding decontextualized subclaims. We decompose and decontextualize sentences from the original Wikipedia page, either from the lead (in the LEAD task) or body (in the BODY task), and provide the lead paragraph (LEAD) or additionally the body paragraph from which the claim originates (BODY) as context for decontextualization. During the pilot annotation, annotators are able to toggle between the subclaim and its decontextualized version to select evidence sentences that together provide the strongest evidence (either supporting or refuting) for the subclaim. Finally, after identifying evidence, annotators determine a support score for the subclaim given that evidence. The bulk annotation provides only the decontextualized subclaim as lead or body claim. Following Wanner et al., we use GPT-4o-mini (OpenAI, 2024) to perform DnD.
@@ A.3 Annotation Interface
  • The subclaim is unfaithful to the meaning of the original sentence (Bad Decontextualization)
@@ A.4 Prompts and Hyperparameters
The prompt used for bulk annotation with GPT-4o-mini is shown in Figure 5 through Figure 9 (divided over multiple pages due to the length of the instructions). This prompt was selected based on highest agreement with the human pilot annotations after numerous manual iterations on other prompts. We used gpt-4o-mini-2024-07-18, the most recent version of the model available. Annotations were generated with temperature 0, with a limit of 2K output tokens to accommodate source texts of up to 126K tokens. Source texts exceeding this limit were truncated, though this was rarely required.
Appendix B Experimental Details and Additional Results
@@ B.1 Qrels for ENTITY
B.2 Fine-Grained Reranking Results
Table 4 shows the BM25 and Rank1 results from Table 2 broken down by number of evidence sentences in the gold annotations for each query (note: R@10 results are omitted, as they are unchanged by reranking the top-10 sentences). These results convincingly demonstrate that the gains brought by leveraging a reasoning model (Rank1) for reranking are not limited to the “easy” cases of single-sentence evidence sets but robustly extend to multi-sentence evidence sets as well.
Task #Sents Model NDCG@5 R@5
LEAD 1 BM25 76.14 86.64
Rank1 84.49 90.67
2 BM25 53.11 47.28
Rank1 61.32 63.47
3 BM25 25.34 27.40
Rank1 35.40 35.28
BODY 1 BM25 75.78 85.71
Rank1 85.72 90.71
2 BM25 59.71 66.13
Rank1 71.90 75.38
3 BM25 45.98 48.28
Rank1 58.10 58.60
Table 4: Gains from reranking the top-10 BM25 evidence sentences for LEAD and BODY using Rank1, broken down by number of evidence sentences for the query/claim. Rank1 shows large gains in all cases.
B.3 Evidence Propagation
§3 briefly presents some analysis on support for the body evidence for a given lead claim. There, we say that we compute an evidence score for a body sentence by taking either the mean or the product of the support scores for its constituent claims (clipping scores $<0$ to 0 in the latter case). We can then compute an overall evidence score for an evidence set by taking the mean or product of the per-sentence scores. Figure 4 plots distributions of overall evidence scores in the PeopleProfiles dev split when applying both mean (blue) and product (orange) aggregation over claims and then (in both cases) applying mean aggregation over sentences. In both cases, we find obtain very middling overall evidence scores—an average of 0.41 for mean and an average of just 0.12 for product.
Figure 4: Distribution of overall evidence scores for PeopleProfiles dev split body evidence with mean- (blue) and product-based (orange) aggregation of body claim support scores for each evidence sentence. See Appendix B.3.
B.4 Retrieval Model Details
@@
The prompts for reranking evidence with Rank1-7B are provided in Figure 10 and Figure 11. Outputs were generated with temperature 0. Context size was set to 16K tokens, with a maximum of 8192 output tokens.
B.5 Use of AI Assistants
@@
Appendix C Further Discussion of Related Work
In §1, we note that the resource most similar to PeopleProfiles is the WiCE dataset from Kamoi et al. (2023), a textual entailment dataset using text-citation pairs from Wikipedia. Here, we discuss some of the key differences between our PeopleProfiles and WiCE, summarized in Table 5. First, support scores in PeopleProfiles are scalar, rather than categorical (SUPPORTED, PARTIALLY-SUPPORTED, NOT-SUPPORTED), as in WiCE, which enables finer-grained analysis of partial support (see §3). Furthermore, PeopleProfiles includes article-internal annotations of claim support (LEAD) in addition to article-external annotations (BODY), whereas WiCE contains only the latter. To our knowledge, ours is the first work to have both types of claim support annotations. We also annotate all lead sentences and all body sentences with attached citations, with WiCE opting to annotate only the SIDE subset (Petroni et al., 2022), containing citations unlikely to support the claim. Although PeopleProfiles annotations are automated by an LLM rather than curated by humans, this allows us to have a dataset over 20×20$\times$20 × as large as WiCE.

### v2

@@ 5 Evidence Retrieval
    LEAD: Retrieve all body evidence sentences for a given lead claim
@@
    BODY: Retrieve all evidence sentences from a single cited source for a given body claim
We also consider a third, entity-level retrieval task:
@@
We report first-stage retrieval results on the PeopleProfiles test set with several widely used models: BM25 (sparse, lexical; Robertson et al., 1995), ColBERTv2 (dense, multi-vector; Khattab and Zaharia, 2020; Santhanam et al., 2022), and Stella-v5-1.5B (dense, single-vector; Zhang et al., 2024). We additionally report reranking results on BM25 outputs with four recent rerankers, discussed later in this section. For LEAD and BODY, we report recall@{5,10} and NDCG@5. For ENTITY, we instead report recall@100, as well as NDCG@{5,100}, since there are $\gg$ 3 evidence sentences per query.
Task Model N@5 R@5 R@10
LEAD ColBERTv2 52.59 57.90 68.18
[…]
BM25 49.90 55.98 65.89
RankZephyr-7B 62.56 62.67 65.89
Rank1-7B 60.55 63.25 65.89
Rank-K-7B 63.33 63.84 65.89
ReasonRank-7B 64.24 64.49 65.89
ReasonRank-32B 64.24 64.41 65.89
BODY ColBERTv2 70.02 76.37 87.16
[…]
BM25 61.52 67.87 78.73
RankZephyr-7B 73.07 74.40 78.73
Rank1-7B 73.02 76.18 78.73
Rank-K-32B 75.01 76.28 78.73
ReasonRank-7B 75.92 76.96 78.73
ReasonRank-32B 76.16 77.04 78.73
Table 3: Evidence retrieval results for the LEAD and BODY tasks. White rows are first-stage retrieval results (“” indicates best results). Blue rows are reranking results on the top 10 evidence sentences from BM25 (best results are bolded). N=NDCG; R=Recall.
5.1 First Stage Retrieval
The first three rows of Table 3 report first stage retrieval results for the LEAD and BODY tasks. On both, we obtain our best results with Stella, which shows 7+ point gains over BM25 across metrics on LEAD and 11+ point gains on BODY. ColBERT also achieves notable improvements over BM25, albeit not quite as large (2+ on LEAD and 8+ on BODY).
The first three rows of Table 4 show results on the ENTITY task, where both Stella and ColBERT again show large improvements over BM25 (up to +10 on NDCG@100; nearly +8 on recall@100), although Stella is no longer the clear winner. ColBERT achieves best results on NDCG@5 (24.53 vs. 23.79), whereas Stella achieves a slight edge on NDCG@100 (40.60 vs. 39.65) and best results on recall@100 (64.37 vs. 62.52).
Collectively, these experiments offer some preliminary evidence that effective evidence retrieval requires more than mere lexical match. While first-stage NDCG scores with dense methods (ColBERT, Stella) are decent—at least for LEAD and BODY—the next section considers how much these results may be improved by leveraging yet more sophisticated reasoning methods for reranking.
Model N@5 N@100 R@100
[…]
BM25 14.59 30.08 56.67
RankZephyr-7B 24.91 35.82 56.67
Rank1-7B 20.54 34.28 56.67
Rank-K-32B 26.98 37.97 56.67
ReasonRank-7B 24.55 36.83 56.67
ReasonRank-32B 26.35 37.60 56.67
Table 4: Evidence retrieval results for the ENTITY task. Blue rows are reranking results on the top 100 evidence sentences from BM25. See Table 3 for further details.
5.2 Reranking via Reasoning
Recent work has shown that reasoning rerankers, which use reasoning chains to produce relevance judgments, achieve substantial gains on other complex retrieval tasks (Weller et al., 2025; Shao et al., 2025; Zhuang et al., 2025, i.a.). Our tasks—particularly the claim-centric ones (LEAD and BODY)—clearly belong in this category, as evidence may be multi-premise, with premises dispersed across the body of the article or source document. Accordingly, we evaluate several such rerankers on all three tasks:
  1. 1.
    Rank1 (7B; Weller et al., 2025): A pointwise reranker based on Qwen 2.5 7B (Qwen et al., 2025) distilled from 635K reasoning traces for MS MARCO (Nguyen et al., 2016) relevance judgments produced by DeepSeek R1 (Guo et al., 2025).
  2. 2.
    Rank-K (32B; Yang et al., 2025): A listwise reranker based on QwQ 32B (Team, 2025) distilled from 50K reasoning traces from R1 on MS MARCO.
  3. 3.
    ReasonRank (7B, 32B; Liu et al., 2025): Listwise rerankers based on Qwen 2.5 instruct models (Qwen et al., 2025) distilled from R1 reasoning traces on relevance judgments across diverse domains (web search, complex QA, coding, math).
[…]
Results on LEAD and BODY are shown in the blue rows of Table 3, where we find large gains in NDCG@5 (10+ points) and recall@5 (6+ points) over BM25 on both tasks across all rerankers. Notably, we achieve our highest scores on LEAD and BODY with the listwise reasoning rerankers—Rank-K and ReasonRank—with the latter achieving best results. Interestingly, however, we do not find substantial improvements in moving from the 7B ReasonRank model to the 32B one.
Broadly, these results echo previous findings that demonstrate superior performance from listwise reasoning rerankers relative to pointwise (Rank1) and non-reasoning models (RankZephyr) on other reasoning-intensive tasks (Yang et al., 2025; Liu et al., 2025). Further, contra Weller et al. (2025), Rank1 generally does not improve upon RankZephyr for our tasks, suggesting that both using listwise reranking and leveraging reasoning are key to strong performance.
5.3 Reasoning Rerankers and Evidence Complexity
A tempting hypothesis is that the effectiveness of listwise reasoning rerankers derives from their comparative advantage in handling more complex evidence—here, examples with multi-premise evidence sets. We explore this in Table 5, which reports results on LEAD broken down by the size of the gold evidence set (# premises), including first-stage BM25 results and reranking results with ReasonRank-7B, Rank1 (as a pointwise baseline), and RankZephyr (as a non-reasoning baseline).
The story these results tell is more complicated than the hypothesis suggests. Although we do find that ReasonRank achieves larger gains over BM25 for 2-premise (+18.3 N@5, +12.4 R@5) and 3-premise (+12.8 N@5, +9.1 R@5) evidence sets than for 1-premise sets (+12.6 N@5, +4.8 R@5), the same is also true of RankZephyr (which does not use reasoning) and Rank1 (which is not listwise). Rather, since ReasonRank achieves the best results among models over all evidence set sizes, a better explanation is that this model simply exhibits the strongest ability to reason about evidential support full stop, irrespective of evidence complexity.
Lastly, we note the large drops in performance observed across all models as one moves from 1-, to 2-, to 3-premise evidence sets. This suggests that there remains substantial room for better cross-premise reasoning, even among the sophisticated listwise rerankers studied here.
5.4 Retrieval Difficulty by Task
@@
A final notable observation about our retrieval results is that retrieval difficulty varies widely across tasks. We obtain highest N@5 scores on BODY, followed by LEAD and then by ENTITY. Intriguingly, this ranking tracks the granularity of the queries, where body claims (BODY) tend to provide the most detailed information, lead claims (LEAD) present key high-level facts, and entity-level queries (ENTITY) are most general, inquiring about the life of the target entity in generic terms. Intuitively, the detailed claims appearing in Wikipedia body sections tend to bear greater lexical and semantic similarity to their supporting source material than the higher-level queries of LEAD or ENTITY do to theirs. At least for BODY and ENTITY, an initial passage-level retrieval stage could plausibly yield improved recall, but the core challenge of reasoning over combinations of fine-grained (sentence-level) premises would still remain.
#Sents Model N@5 R@5 R@10
1 BM25 76.13 86.60 91.78
RankZephyr-7B 86.53 89.64 91.78
Rank1-7B 84.49 90.67 91.78
ReasonRank-7B 88.68 91.38 91.78
2 BM25 47.29 53.18 66.51
RankZephyr-7B 63.90 63.28 66.51
Rank1-7B 61.32 63.47 66.51
ReasonRank-7B 65.62 65.57 66.51
3 BM25 25.25 27.26 39.11
RankZephyr-7B 37.01 34.88 39.11
Rank1-7B 35.40 35.28 39.11
ReasonRank-7B 38.08 36.37 39.11
@@
Table 5: Select retrieval and reranking results on LEAD broken down by number of gold evidence sentences.
@@ 6 Conclusion
We have presented a study of evidential support and retrieval on Wikipedia and have introduced PeopleProfiles, a large dataset of fine-grained, multi-level evidential support annotations on nearly 1,500 Wikipedia articles and their cited sources. We have shown that: (1) a sizable fraction of Wikipedia lead claims are unsupported by their body sections, and Wikipedia body claims by their (publicly accessible) cited sources; (2) evidence retrieval for these claims grows much more challenging as the generality of the queries (LEAD $\rightarrow$ BODY $\rightarrow$ ENTITY) and the evidence complexity (1 $\rightarrow$ 2 $\rightarrow$ 3 premises) increases; and lastly (3) new reasoning-based rerankers enable much more effective retrieval of complex evidence relative to traditional (sparse or dense) methods. We release PeopleProfiles to aid future work on claim verification and on advancing understanding of Wikipedia as a knowledge source for modern NLP.
@@
We acknowledge several limitations of our work. First, PeopleProfiles focuses only on Wikipedia articles about people. We chose this focus because biographies tend to have a higher proportion of uncontroversial facts relative to other domains (e.g. concepts or events) and because multiple prior works in this area also focus on people (Min et al., 2023; Jiang et al., 2024; Gunjal and Durrett, 2024). However, it is possible support distributions or evidence retrieval difficulty could differ in other domains. (Granted, we believe domain shift would likely strengthen many of the claims we make, as discussed in §3.3.)
Second, as we emphasize in the main text, our claims about evidential support extend only to publicly accessible, digital sources. We therefore cannot make conclusions about support across all source types in Wikipedia. The number of paid licenses and the level of access to print resources required for this are simply infeasible to obtain.
Third, we leverage GPT-4o-mini as an annotator to facilitate our large-scale bulk data collection. While the agreement we observe between this model and our human annotations is strong (§3), LLMs have their own response biases and may not be fully calibrated when providing scalar judgments (Lovering et al., 2024).
Finally, as we note in §4.3, real-world citation practices can differ from what is espoused in documented standards—where citations often do not appear on a sentence that would seem to require one, but where they may appear elsewhere in the same paragraph. Since we restrict our annotations to body sentences that do bear citations, we thus cannot draw conclusions about sentences whose source may appear as a citation on some other sentence. However, we suspect that our estimate of the proportion of supported body claims decomposed from citation-bearing sentences is an upper bound on the proportion we would obtain for all citation-worthy sentences.
@@
PeopleProfiles’s use of sources from MegaWika 2.0 and our release of this data (via a CC-BY-4.0-SA license) is consistent with MegaWika 2’s own CC-BY-4.0-SA license. Our principle transformation of the original Wikipedia articles consists in the claim decomposition, which is performed by an LLM (GPT-4o-mini), and which can occasionally result in (sub)claims that misrepresent the article’s original content and thus (potentially) facts about the subject. Although our claim decompositions are generally very faithful to the original texts, users of PeopleProfiles should be aware of this possibility.
@@ A.1 Annotator Demographics
The first three authors of this work—all native English-speaking graduate students in NLP or professional NLP researchers—conducted the human pilot annotations. These authors also jointly produced the annotation instructions beforehand. None was compensated beyond their co-authorship on this work.
@@ A.2 Claim Decomposition
We use the “DnD” decomposition and decontextualization method introduced by Wanner et al. (2024b), which uses an LLM prompt-based method for obtaining subclaim decompositions and the corresponding decontextualized subclaims. We decompose and decontextualize sentences from the original Wikipedia page, either from the lead (in the LEAD task) or body (in the BODY task), and provide the lead paragraph (LEAD) or additionally the body paragraph from which the claim originates (BODY) as context for decontextualization. During the pilot annotation, annotators are able to toggle between the subclaim and its decontextualized version to select up to three sentences that together provide the strongest evidence (either supporting or refuting) for the subclaim. Finally, after identifying evidence, annotators determine a support score for the subclaim given that evidence. The bulk annotation provides only the decontextualized subclaim in the prompt. Following Wanner et al., we use GPT-4o-mini (OpenAI, 2024) to perform DnD.
@@ A.3 Annotation Interface
  • The subclaim is unfaithful to the meaning of the sentence from which it was decomposed (Bad Decontextualization)
@@ A.4 Prompts and Hyperparameters
The prompt used for bulk annotation with GPT-4o-mini is shown in Figure 6 through Figure 10 (divided over multiple pages due to the length of the instructions). This prompt was selected based on highest agreement with the human pilot annotations after numerous manual iterations on other prompts. We used gpt-4o-mini-2024-07-18, the most recent version of the model available at the time. Annotations were generated with temperature 0, with a limit of 2K output tokens to accommodate source texts of up to 126K tokens. Source texts exceeding this limit were truncated, though this was rarely required.
Appendix B Experimental Details
@@ B.1 Qrels for ENTITY
B.2 Retrieval Model Details
@@
Rank1 outputs were generated with temperature 0 and a maximum of 8,192 output tokens. We adopted the generation temperatures for listwise rerankers from the defaults listed in their public repositories and limited their outputs to 8K tokens. For listwise reranking, we used a window size of 20 with stride 10.
B.3 Reranking Prompts
Figure 11, Figure 12, Figure 13, and Figure 14 list the task-specific modifications made to the prompts used in the original reranker implementations.
B.4 Use of AI Assistants
@@
Appendix C Additional Results
C.1 Additional Reranking Results
Table 6 shows select reranking results ($k=10$) on top of Stella-1.5B-v5 outputs (rather than BM25, as in the main text) for the LEAD and BODY tasks. Similar to the results reported in Table 3, we obtain the highest scores with ReasonRank on both tasks, with comparably large gains over first-stage retrieval.
Table 7 shows the same results for the ENTITY task ($k=100$). Here again, ReasonRank achieves the best scores, although the gains relative to first-stage retrieval are much smaller compared to those based on BM25 outputs (Table 4).
Task Model N@5 R@5 R@10
LEAD Stella-1.5B-v5 59.67 64.92 75.12
RankZephyr-7B 68.96 69.92 75.12
Rank-K-32B 70.0 71.44 75.12
ReasonRank-7B 71.11 72.36 75.12
BODY Stella-1.5B-v5 74.27 80.69 90.41
RankZephyr-7B 81.28 83.88 90.41
Rank-K-32B 83.54 86.40 90.41
ReasonRank-7B 84.31 87.22 90.41
Table 6: Reranking results for select models on outputs from Stella-1.5B-v5 on the LEAD and BODY tasks (best results are bolded). N=NDCG; R=Recall.
Model N@5 N@100 R@100
Stella-1.5B-v5 23.79 40.60 64.37
RankZephyr-7B 25.28 41.14 64.37
Rank-K-32B 25.47 42.08 64.37
ReasonRank-7B 25.99 42.54 64.37
Table 7: Reranking results for select models on outputs from Stella-1.5B-v5 on the ENTITY task (best results are bolded). N=NDCG; R=Recall.
Appendix D WiCE Comparison
Table 8 summarizes key differences between our PeopleProfiles and the most similar existing benchmark, WiCE (Kamoi et al., 2023). See §2 for a detailed comparison between PeopleProfiles, WiCE, and other related works.